OpenStudy (anonymous):

If I give someone an problem can you tell me if the answer I got is correct?

OpenStudy (anonymous):

That depends.

OpenStudy (anonymous):

Haha @ElmerHigglesworth true true...

OpenStudy (anonymous):



OpenStudy (anonymous):

I got x= -6

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

Hmm... this is a toughie... might take me a second.

OpenStudy (anonymous):

Or -1/6

OpenStudy (anonymous):

I am thinking -1/6

OpenStudy (anonymous):

\[\eqalign{ &\frac{5}{4x}+\frac{2}{1}=\frac{1}{x} \\ &\frac{8x+6}{4x}=\frac{1}{x} \\ &8x^2+6x=4x \\ &8x^2-2x=0 \\ &2(4x^2-x)=0 \\ &2x(4x-1)=0 \\ }\]

OpenStudy (anonymous):

I got \[x∈{{2/5,(−2)}}\]

OpenStudy (anonymous):

So would that be NO SOLUTION?

OpenStudy (anonymous):

If the answer was \(-\frac{1}{6}\), then: \[\eqalign{ &\left.\left(\frac{5}{4x}+\frac{2}{1}=\frac{1}{x}\right)\right|_{x=-\frac{1}{6}} \\ &\left(\frac{5}{4\times-\frac{1}{6}}+\frac{2}{1}=\frac{1}{-\frac{1}{6}}\right) \\ }\]

OpenStudy (anonymous):

Yeah, pretty much no solution.

OpenStudy (anonymous):

Err wait yeah... The solution to: \[2x(4x+1)=0\] Is \((x=0)\phantom{.}or\phantom{.}(x=-\frac{1}{4})\)

OpenStudy (anonymous):

\[\frac{ 1 }{ x } + \frac{ 2 }{ 3x } = \frac{ 1 }{ 3}\]

OpenStudy (anonymous):

So how would I know what to put for x?

OpenStudy (anonymous):

Wait. If this is another equation for another question, put it under a separate question and tag people to answer it. Don't put more than one question on one space.